The wooden box is thrown with some initial speed in the horizontal plane and after crossing the twilight 5 m for 2 s it stops. Calculate the coefficient of sliding friction.
1
Expert's answer
2020-12-15T11:47:44-0500

Let's first find the initial velocity of the box:


"d=\dfrac{v_0+v}{2}t,""v_0=\dfrac{2d}{t}-v=\dfrac{2\cdot 5\ m}{2\ s}-0=5\ \dfrac{m}{s}."

Then, we can find the deceleration of the box:


"v=v_0 + at,""a=\dfrac{v-v_0}{t}=\dfrac{0\ \dfrac{m}{s}-5\ \dfrac{m}{s}}{2\ s}=-2.5\ \dfrac{m}{s^2}."

The sign minus means that the box decelerates.

Finally, from the Newton's Second Law of Motion we can find the coefficient of sliding friction:


"\sum F=ma,""F_{k.fr.}=ma,""-\mu_k mg=ma,""\mu_k=-\dfrac{a}{g}=-\dfrac{-2.5\ \dfrac{m}{s^2}}{9.8\ \dfrac{m}{s^2}}=0.25"

Answer:

"\mu_k=0.25"